OpenStudy (anonymous):

19u+4v-3u-9v= Collect like terms

OpenStudy (anonymous):

16u - 5v

OpenStudy (anonymous):

I took 19u-3u=16 4v-9v=-5 I am not sure what to do from here

OpenStudy (anonymous):

where'd you get 16 and 5 from?

OpenStudy (anonymous):

well I 19 minus 3

OpenStudy (anonymous):

you put 19u + 4v -3u - 9v = combine the u terms and the v terms. this simplifies to 16u - 5v

OpenStudy (anonymous):

right

OpenStudy (anonymous):

which is 11

OpenStudy (anonymous):

that isn't 11. u and v are different variables, so you can't combine them.

OpenStudy (anonymous):

16u - 5v is the final answer.

OpenStudy (anonymous):

ok thanks a lot I was trying to remember how to do it cause I have done it.

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

can you please look at the other problem to if i post it here

OpenStudy (anonymous):

yea post it.

OpenStudy (anonymous):

Train A and B are traveling in the same direction on parallel tracks. Train A is traveling at 60 mph and b is traveling 80 mph. Train A passes at 7:25pm. Train B passes the same station at 7:37pm. What time will train B catch up to train A?